OpenStudy (anonymous):

A four-unit class expects each student to dedicate 10 hours a week to studying for the course. How many hours should a student expect to study for a particular class that is worth six units?

OpenStudy (bibby):

Set up a proportion\[\huge \frac{ 4unit }{ 10hours }= \frac{ ? }{ ? }\] what goes on the right side?

OpenStudy (gorv):

first find no of hour dedicated to one unit that is = no of hoers dedicated / no of units no of hour dedicated to one unit =10/4=5/2 for 6 unit multiply it with 6 =(5/2)*6=15 hr/week @brookelovesyewh